OpenStudy (anonymous):

how to factor this 2(x^3)-5

OpenStudy (anonymous):

\[ ( x \sqrt[3]{2})^3 - (\sqrt[3]{5})^3\]=

OpenStudy (anonymous):

now \[a^3 -b^3 = (a-b)(a^2 +ab +b^2)\]

OpenStudy (anonymous):

where \[a= x \sqrt[3]{2}\] \[b= \sqrt[3]{5}\]